OpenStudy (anonymous):

Find the value of the derivative of f(θ) = 2 sin θ − θ at (0,0)

OpenStudy (anonymous):

f(x) = 2 sin x - x f'(x) = 2 cos x -1 At ( 0 , 0 ) , x-coordinate = 0 , f'(x) = 2 cos 0 - 1 = 1

OpenStudy (anonymous):

if you dont understand some parts just tell :D

OpenStudy (anonymous):

!o i got it! Thank you! I saw the θ and freaked for a second lmao

OpenStudy (anonymous):

GL :))